Steven Company has fixed costs of $185,484. The unit selling price, variable cost per unit, and contribution margin per unit for the company's two products are provided below.

Product Selling Price per Unit Variable Cost per Unit Contribution Margin per Unit
X $768 $288 $480
Y 323 173 150

The sales mix for Products X and Y is 60% and 40%, respectively. Determine the break-even point in units of X and Y. Round answers to the nearest whole number.
units of X____
units of Y____
Answer #1

Weighted average contribution margin=Respective contribution margin*Respective sales mix

=(480*0.6)+(150*0.4)=348

Breakeven=Fixed cost/Weighted average contribution margin

=185,484/348

=533 units

Hence breakeven point of:

X=533*60%=320 units

Y=533*40%=213 units.